The package easytable

Enrico Bertolazzi

Laboratorio di Matematica Applicata e Meccanica Computazionale

Dipartimento di Ingegneria Meccanica e Strutturale Università degli Studi di Trento Mesiano, Trento, Italy

           mailto:enrico.bertolazzi@ing.unitn.it

Abstract—The easytable package is a macro package for writing tables, with equal column widths or equal rows heights or both, with various kinds of rules (lines) between rows and columns. It uses an array/tabular-like syntax.

1 How to use it

Load the package in the usual way:

\documentclass{article}
.
.
\usepackage[thinlines,thiklines]{easytab}
.
.

The options thinlines, and thiklines are self explanatory. easytab provides the tab environment which is a simple re-implementation of the array/tabular environment, with some limitation and some additional features. The syntax is

\begin{TAB}`(eq)'`[ex]'`{cc...c}'`{cc...c}'
   a & b & ... & n \\
.
\end{TAB}
or
\begin{TAB}`(eq,mx,my)'`[ex,MX,MY]'`{cc...c}'`{cc...c}'
   a & b & ... & n \\
.
\end{TAB}

• (eq) or (eq,mx,my). By eq you can balance the rows or the column or both, as shown in this table:

Table 1.

value of x effect
@  no balancing
r  equal rows heights
c  equal column widths
b  equal rows heights and equal column widths
e  equal rows heights and column widths

By mx and my you can modify the minimum size of the box in the TAB environment. This must be a valid measure e.g. 2pt. This is useful in writing matrices an vectors.

• [ex] or [ex,MX,MY]. By “ex” you can specify the amount of extra space around the item in the TAB environment. The default is 2pt. By MX and MY you can modify the minimum size of the whole table in the TAB environment. This must be a valid measure e.g.10cm.

• The first `{cc...c}' is the definition of the columns and their alignment. The possible alignment for the columns are:

Table 2.

c centering
l flush left
r flush right

• The second `{cc...c}' is the definition of the rows their alignment. The possible alignment for the rows are:

Table 3.

c centering
t flush top
b flush bottom

IMPORTANT: The package can manage matrices with a maximum of 30 rows by 30 columns. IMPORTANT: The functionality of the environment “TAB” is the same of the environment “BMAT” the only difference is that its entries are in “text” mode not in “mathematic” mode. Please read the documentation of the package easybmat to understand how to use “TAB”.